# Panoramic Out-of-Distribution Segmentation

### v1

@@ 3.4 Bilevel Prompt Distribution Learning
In-distribution learning. To prevent the model from misinterpreting intra-class pixel embeddings when distinguishing out-of-distribution from in-distribution pixel embeddings $\{m_{i}{\mid}i{=}1,\dots,N_{i}\}{\in}\mathbb{R}^{N_{i}\times C_{m}}$, thereby improving closed-set segmentation performance, we introduce an in-distribution loss $\mathcal{L}_{pixel}$. As shown in Fig. 7, it guides the model’s learning process by enhancing intra-class pixel embedding compactness while preserving inter-class separability.
First, we minimize the Euclidean distance between each pixel embedding and the corresponding class center, optimizing intra-class compactness, which can be represented as follows:
@@
Figure 7: In-distribution Learning. Enhancing intra-class compactness: the similarity of pixel embeddings within the same class is reinforced, resulting in a more compact distribution of intra-class features. Improving inter-class separation: the embedding distance between different classes is expanded, thereby strengthening the discriminability of inter-class features.
@@
$\mathcal{L}_{ori}=\frac{1}{K+1}\sum_{i=0}^{K}\sum_{o\neq i,n}\omega_{o}(1+\cos(\vec{t}_{n},\leavevmode\nobreak\ \vec{t}_{o})),$ (11)
@@
Figure 8: Distribution Learning. Isolation of out-of-distribution pixels: A distinct boundary is established in the feature space to separate out-of-distribution regions from known class regions, thereby reducing the model’s overconfidence in out-of-distribution areas. Aggregation of out-of-distribution pixels: through semantic guidance, out-of-distribution pixel embeddings are clustered into specific semantic spaces, preventing their dispersion and interference with the feature distribution of known classes.
Distribution learning. Building upon in-distribution learning, we propose a higher-level distribution learning mechanism to enhance the separability of out-of-distribution embeddings $O_{i}$ and achieve a more precise delineation of the overall feature space.
To establish a robust foundation for semantic alignment in subsequent distribution region partitioning and enable the model to more accurately capture the feature distribution patterns of known classes, we first minimize the distance between class text prompts $\{T_{i}\mid i=0,\dots,K\}$ and in-distribution prompts $P_{in}$ in the feature space. The in-distribution compactness loss $\mathcal{L}_{ind}$ promotes that the feature distribution of known classes aligns with the in-distribution semantic information:
@@
$\mathcal{L}_{inDir}=\frac{1}{K+1}\sum_{i=0}^{K}1-\cos(\vec{d}_{in},\leavevmode\nobreak\ \vec{d}_{out}).$ (17)
@@
$\mathcal{L}_{inlier}=\alpha\leavevmode\nobreak\ \mathcal{L}_{ind}+\mathcal{L}_{inDir},$ (18)
@@
Furthermore, to enhance the model’s ability to identify out-of-distribution regions, we employ semantic guidance to cluster out-of-distribution pixel embeddings $\{O_{1},\dots,O_{N_{o}}\}$ into specific regions of the semantic space. This approach not only preserves the stability of the known class feature distribution but also provides clear semantic attribution for out-of-distribution features. Additionally, it distinctly delineates out-of-distribution regions in the feature space, establishing a clear boundary with known class regions. As a result, it reduces the model’s overconfidence in out-of-distribution areas and enhances the reliability of detection.
@@
where $\{D_{i,out},\leavevmode\nobreak\ D_{i,in}\}=\{\|O_{i}-P_{out}\|_{2}^{2},\leavevmode\nobreak\ \|O_{i}-P_{in}\|_{2}^{2}\}$, $d$ quantifies the degree of difference between $P_{in}$ and $P_{out}$ within the semantic space. The distribution learning loss $\mathcal{L}_{distri}$ is formulated as a sum of the in-distribution loss $\mathcal{L}_{inlier}$ and the out-of-distribution pixel loss $\mathcal{L}_{outlier}$, which is calculated as follows:
@@ 3.5 Loss Settings
$\mathcal{L}_{mask}=\lambda_{bce}\leavevmode\nobreak\ \mathcal{L}_{bce}+\lambda_{dice}\leavevmode\nobreak\ \mathcal{L}_{dice},$ (22)
@@
$\mathcal{L}=\mathcal{L}_{mask}+\lambda_{cls}\leavevmode\nobreak\ \mathcal{L}_{cls}.$ (23)
@@
$\mathcal{L}_{oe}=\mathcal{L}+\mathcal{L}_{rba}+\lambda_{bpdl}\leavevmode\nobreak\ \mathcal{L}_{bpdl},$ (24)
@@
4 Established Benchmarks
Figure 9: Distribution details of DenseOoS. DenseOoS contains $30$ distinct out-of-distribution classes, divided into $5$ main categories: Animals, Garbage, Furniture, Vehicles & Tools, and Others.
4.1 Overview of the Benchmarks
@@
In this work, we establish two novel benchmarks: DenseOoS and QuadOoS, specifically designed for PanOoS. $360^{\circ}$ panoramas have a broader FoV and many small objects in the images, which exaggerate the costs of creating pixel-wise annotations in unconstrained surroundings. From the perspective of the task, we aim to overcome the scarcity of PanOoS testbeds and effectively evaluate the impact of mitigating the background clutter and severe distortions on out-of-distribution segmentation performance when unmasking the narrow FoV. Furthermore, to explore PanOoS in real-world, unconstrained, unstructured, and dense-motion scenarios, we utilize a quadruped mobile robot to capture images in outdoor motion scenes with outliers, creating a real-world benchmark. Our objective with two benchmarks is to provide a comprehensive evaluation of methods capable of performing PanOoS. For comparison to previous pinhole-OoS methods, we adapt Area under PRC Curve (AuPRC/AP), and False Positive Rate at a true positive rate of $95\%$ (FPR95) for panoramic out-of-distribution segmentation pixel-wise evaluation, mean Intersection-over-Union (mIoU) for closed-set segmentation evaluation. We unfold a comprehensive explanation in Sec 5.3.
4.2 DenseOoS
Figure 10: Examples from the DenseOoS dataset. The background of the images is highly cluttered, and the distortions are significant. Outliers are distributed in reasonable areas within the image.
[…]
The overall process of dataset generation and construction is as follows. First, all text prompts for image inpainting (a good photo of a $\{class\}$) are determined based on the DensePASS validation dataset; Next, we select appropriate regions to be inserted (e.g., the road and sidewalk) and the selected rectangular area is scaled to match the resolution of the real panoramic images. Then, we set different random seeds and use the POC algorithm to batch-generate images and outlier annotations. Finally, the suitable images for constructing the DenseOoS dataset are determined through preliminary manual screening, independent screening, and majority voting. Specifically, a total of $14,031$ images are generated, of which $1,244$ are retained after preliminary manual screening. Following an independent screening and voting process by three examiners, we ultimately select $1,000$ images for the construction of DenseOoS. As shown in Fig. 10, the outliers are diverse and distributed within reasonable areas of the images. A significant portion of these outliers extends beyond the pinhole narrow FoV, making them particularly suitable for the evaluation of the PanOoS task.
4.3 QuadOoS
(a) Data collection platform for QuadOoS dataset.
Frames Resolution Locations Objects Annotations
Test 106 $2048{\times}400$ 8 $\geq$18 735
(b) Details of the QuadOoS dataset.
(c) Y-axis pixel positions of motion trajectories.
To further explore the challenges of PanOoS task in real applications, we collect panoramic images featuring outliers in unconstrained, unstructured, and dense-motion real scenes, manually annotating these outliers to establish the first Quadruped Panoramic Out-of-distribution Segmentation dataset, QuadOoS. This dataset provides a crucial foundation for panoramic out-of-distribution segmentation research, facilitating the evaluation and optimization of PanOoS models in complex environments.
We develop a mobile platform using a quadruped robot, and the robot was selected for its biomimetic gait, a complex form of locomotion found in nature, which introduces additional challenges for PanOoS. The robot measures $70cm{\times}31cm{\times}40cm$, with a maximum payload capacity of $7kg$. It can navigate vertical obstacles up to $15cm$ and inclines up to $30^{\circ}$, making it highly maneuverable in everyday environments. With $12$ joint motors, the robot replicates realistic walking motions at speeds up to $2.5m/s$. For sensing, we use a customized panoramic annular camera to capture a broader FoV ($360^{\circ}{\times}70^{\circ}$). The camera has a pixel size of $3.45{\mu}m{\times}3.45{\mu}m$, a resolution of $5$ million effective pixels. Mounted on the quadruped robot (see Fig. 11(a)), the camera ensures an unobstructed, optimal field of view.
Using this platform, we conduct data collection in campus environments with outliers, creating a Quadruped Panoramic Out-of-distribution Segmentation (QuadOoS) dataset. In particular, we collect data across $8$ blocks in and around the campus, with each block averaging approximately $50m$ in length. We pre-place the outliers and subsequently perform the actual data collection using the platform outlined in the main text. The collected images consist of a total of $4,800$ frames ($600$ frames per block), each with an original size of $2048{\times}476$ pixels. To avoid misleading the model due to excessive exposure at the center of the lens and to maintain consistency with DenseOoS, we crop the upper portion of the original images (i.e., the sky), which does not affect fair evaluation. Subsequently, to ensure the diversity of outliers and the satisfactory imaging quality, we retain a final selection of $106$ original panoramic images after manual screening by five examiners, thereby constructing the QuadOoS dataset. Detailed information is provided in Fig. 11(b), and see Appendix B.1 for details of the types of outliers.
@@ 4.3 QuadOoS
Table I: Comparison of different OoS datasets. Referring to POC [80], we qualitatively compare datasets on two main axes. We also score them as either good (), medium ((✓)) or bad ().
Figure 12: Examples from the established QuadOoS dataset. The images come with challenges such as uneven exposure, color inconsistencies due to the broader FoV, and increased motion blur. A large number of small outliers are scattered in the background.
5 Experiment Results and Analysis
5.1 Datasets
@@
5.2 Implementation Details
Closed-Set Training. We follow [26] and adopt the same training recipe and losses without any special design. The difference is that we adopt the ConvNeXt-Large CLIP [73] from OpenCLIP [82], pre-trained on the LAION-2B dataset [83], and freeze its weights. In reference to previous research works [25, 24], only one decoder layer with $100$ queries is used in the transformer decoder. Both dimensions $C_{m}$ and $C_{q}$ are set to $256$. The model is trained for $90k$ iterations using a batch size of $16$ on $4$ NVIDIA GeForce RTX 3090 GPUs. The training is optimized with AdamW optimizer and weight decay $0.05$. We employ the initial learning rate $1{\times}10^{-4}$ and a polynomial schedule.
@@ 5.2 Implementation Details
5.3 Evaluation Metrics
@@
5.4 Results of Panoramic Out-of-distribution Segmentation
Method PanOoS QuadOoS Mean
PEBAL [29] 16.39 7.59 46.29 11.34 39.25 13.87 23.42
DenseHybrid [30] 24.44 7.74 45.19 19.09 33.86 21.77 20.80
EAM [23] 17.97 6.46 61.58 33.05 79.66 25.51 43.06
Mask2Anomaly [25] 36.28 8.79 48.10 25.93 92.55 31.10 50.67
RbA [24] 51.31 21.87 57.54 34.12 90.34 42.71 56.11
Ours 85.56 0.45 63.30 48.13 85.22 66.85 42.84
Method mIoU
DAFormer [31] 44.53
Trans4PASS (T) [8] 44.93
Trans4PASS (S) [8] 49.35
Trans4PASS+ (T) [32] 49.29
Trans4PASS+ (S) [32] 50.11
360SFUDA++ [46] 50.46
Ours 63.30
Figure 13: Visual comparisons of PanOoS. We observe that when existing pinhole-OoS methods are transferred to PanOoS, current state-of-the-art methods generally struggle with misidentifying complex background as outliers. Although EAM [23] has a low false positive rate, it incorrectly identifies bicycles as outliers and nearly overlooks the actual outliers. In contrast, Mask2Anomaly [25] and RbA [24] suffer from large false positives in the background region, whereas POS delivers more precise PanOoS across the $360^{\circ}$.
@@ 5.4 Results of Panoramic Out-of-distribution Segmentation
In Fig. 13, we visualize the PanOoS results produced by POS and other mask-transformer-based methods. Notably, POS effectively suppresses false positives in background regions, especially at the boundaries separating inliers, while better preserving the smoothness of the outlier map compared to other methods, despite sharing the same mask classification training paradigm. Benefiting from the generalization ability of vision-language model, Prompt-based Restoration Attention (PRA) and Bilevel Prompt Distribution Learing (BPDL), POS performs excellently on the two benchmarks for PanOoS, perfectly adapting to the background clutter and pixel distortions caused by the broader FoV of panoramic images.
5.5 Results of Pinhole Out-of-distribution Segmentation

### v2

@@ III-D Bilevel Prompt Distribution Learning
In-distribution learning. To prevent the model from misinterpreting intra-class pixel embeddings when distinguishing out-of-distribution from in-distribution pixel embeddings $\{m_{i}{\mid}i{=}1,\dots,N_{i}\}{\in}\mathbb{R}^{N_{i}\times C_{m}}$, thereby improving closed-set segmentation performance, we introduce an in-distribution loss $\mathcal{L}_{pixel}$. As shown in Fig. 7, it guides the model’s learning process by enhancing intra-class pixel embedding compactness while preserving inter-class separability. First, we minimize the Euclidean distance between each pixel embedding and the corresponding class center, optimizing intra-class compactness, which can be represented as follows:
@@
Figure 7: In-distribution Learning. Enhancing intra-class compactness: the similarity of pixel embeddings within the same class is reinforced, resulting in a more compact intra-class distribution. Improving inter-class separation: the embedding distance between different classes is expanded, thereby strengthening the discriminability of inter-class features.
@@
$\mathcal{L}_{ori}=\frac{1}{K+1}\sum_{i=0}^{K}\sum_{o\neq i,n}\omega_{o}(1+\cos(\vec{t}_{n},~\vec{t}_{o})),$ (11)
@@
Distribution learning. Building upon the aforementioned learning, a higher-level distribution learning mechanism is proposed to enhance the separability of out-of-distribution embeddings $O_{i}$ and achieve a more precise delineation of the overall feature space. To establish a robust foundation for semantic alignment in subsequent distribution region partitioning and enable the model to more accurately capture the feature distribution patterns of known classes, we first minimize the distance between class text prompts $\{T_{i}\mid i=0,\dots,K\}$ and in-distribution prompts $P_{in}$ in the feature space. The in-distribution compactness loss $\mathcal{L}_{ind}$ promotes that the feature distribution of known classes aligns with the in-distribution semantic information:
@@
$\mathcal{L}_{inDir}=\frac{1}{K+1}\sum_{i=0}^{K}1-\cos(\vec{d}_{in},~\vec{d}_{out}).$ (17)
@@
$\mathcal{L}_{inlier}=\alpha~\mathcal{L}_{ind}+\mathcal{L}_{inDir},$ (18)
@@
Figure 8: Distribution Learning. Isolation of out-of-distribution pixels: A distinct boundary is established in the feature space to separate out-of-distribution regions from known class regions, thereby reducing the overconfidence for outliers. Aggregation of out-of-distribution pixels: through semantic guidance, out-of-distribution pixel embeddings are clustered into specific semantic spaces, preventing their dispersion and interference with the feature distribution of known classes.
Furthermore, to enhance the model’s ability to identify out-of-distribution regions, we employ semantic guidance to cluster out-of-distribution pixel embeddings $\{O_{1},\dots,O_{N_{o}}\}$ into specific regions of the semantic space. This approach not only preserves the stability of the known class feature distribution but also provides clear semantic attribution for out-of-distribution features. Additionally, it distinctly delineates out-of-distribution regions in the feature space, establishing a clear boundary with known class regions. As a result, it reduces the model’s overconfidence for outliers and enhances the reliability of detection.
@@
where $\{D_{i,out},~D_{i,in}\}=\{\|O_{i}-P_{out}\|_{2}^{2},~\|O_{i}-P_{in}\|_{2}^{2}\}$, $d$ quantifies the degree of difference between $P_{in}$ and $P_{out}$ within the semantic space. The distribution learning loss $\mathcal{L}_{distri}$ is formulated as a sum of the in-distribution loss $\mathcal{L}_{inlier}$ and the out-of-distribution pixel loss $\mathcal{L}_{outlier}$, which is calculated as follows:
@@ III-E Loss Settings
$\mathcal{L}_{mask}=\lambda_{bce}~\mathcal{L}_{bce}+\lambda_{dice}~\mathcal{L}_{dice},$ (22)
@@
$\mathcal{L}=\mathcal{L}_{mask}+\lambda_{cls}~\mathcal{L}_{cls}.$ (23)
@@
$\mathcal{L}_{oe}=\mathcal{L}+\mathcal{L}_{rba}+\lambda_{bpdl}~\mathcal{L}_{bpdl},$ (24)
@@
IV Established Benchmarks
IV-A Overview of the Benchmarks
@@
In this work, we establish two novel benchmarks: DenseOoS and QuadOoS, specifically designed for PanOoS. $360^{\circ}$ panoramas have a broader FoV and many small objects in the images, which exaggerate the costs of creating pixel-wise annotations in unconstrained surroundings. From the perspective of the task, we aim to overcome the scarcity of PanOoS testbeds and effectively evaluate the impact of mitigating the background clutter and severe distortions on out-of-distribution segmentation performance when unmasking the narrow FoV. Furthermore, to explore PanOoS in real-world, unconstrained, unstructured, and dense-motion scenarios, we utilize a quadruped mobile robot to capture images in outdoor motion scenes with outliers, creating a real-world benchmark. Our objective with two benchmarks is to provide a comprehensive evaluation of methods capable of performing PanOoS. For comparison to previous pinhole-OoS methods, we adapt Area under PRC Curve (AuPRC/AP), and False Positive Rate at a true positive rate of $95\%$ (FPR95) for panoramic out-of-distribution segmentation pixel-wise evaluation, mean Intersection-over-Union (mIoU) for closed-set segmentation evaluation. We unfold a comprehensive explanation in Sec V-C.
@@
TABLE I: Comparison of different OoS datasets. Referring to POC [81], we qualitatively compare datasets on two main axes, and score them as good (), medium ((✓)) or bad ().
IV-B DenseOoS
The normal labels cover $19$ classes that align with the Cityscapes [15], while the abnormal labels contain $30$ classes of outliers as demonstrated in Fig. 9. To ensure the high quality and rationality of the generated images, particularly that the outliers are generated in realistic areas within the real street backgrounds, our screening process follows the “preliminary screening $\rightarrow$} independent screening $\rightarrow$} voting” workflow, eliminating unreasonable object locations, and normal situations. Specifically, a total of $14,031$ images are generated, of which $1,244$ are retained after preliminary manual screening. Following an independent screening and voting process by three examiners, we ultimately select $1,000$ images for the construction of DenseOoS. More details are in the Appendix. As shown in Fig. 10, the outliers are diverse and distributed within reasonable areas of the images. A significant portion of these outliers extends beyond the pinhole narrow FoV, making them particularly suitable for the evaluation of the PanOoS task.
IV-C QuadOoS
To further explore the challenges of the PanOoS task in real applications, we collect panoramic images featuring outliers in unconstrained, unstructured, and dense-motion real scenes, manually annotating these outliers to establish the first Quadruped Panoramic Out-of-distribution Segmentation dataset, QuadOoS. This dataset provides a crucial foundation for panoramic out-of-distribution segmentation research, facilitating the evaluation and optimization of PanOoS models in complex environments. We develop a mobile platform using a quadruped robot, and the robot was selected for its biomimetic gait, a complex form of locomotion found in nature, which introduces additional challenges for PanOoS. It can navigate vertical obstacles up to $15cm$ and inclines up to $30^{\circ}$, making it highly maneuverable in everyday environments. With $12$ joint motors, the robot replicates realistic walking motions at speeds up to $2.5m/s$. For sensing, we use a customized panoramic annular camera to capture a broader FoV ($360^{\circ}{\times}70^{\circ}$). The camera has a pixel size of $3.45{\mu}m{\times}3.45{\mu}m$, a resolution of $5$ million effective pixels. Mounted on the quadruped robot (see Fig. 11(a)), the camera ensures an optimal field of view.
Figure 9: Distribution details of DenseOoS. DenseOoS contains $30$ distinct outliers, divided into 55 main categories: Animals, Garbage, Furniture, Vehicles & Tools, and Others.
Figure 10: Examples from the DenseOoS dataset. The background of the images is highly cluttered, and the distortions are significant. Outliers are distributed in reasonable areas within the image. Zoom in for a better view.
Using this platform, we conduct data collection in campus environments with outliers, creating a Quadruped Panoramic Out-of-distribution Segmentation (QuadOoS) dataset. In particular, we collect data across $8$ blocks in and around the campus, with each block averaging approximately $50m$ in length. The collected images consist of a total of $4,800$ frames ($600$ frames per block), each with an original size of $2048{\times}476$ pixels. To avoid misleading the model due to excessive exposure at the center of the lens and to maintain consistency with DenseOoS, we crop the upper portion of the original images (i.e., the sky), which does not affect fair evaluation. Subsequently, to ensure the diversity of outliers and the satisfactory imaging quality, we retain a final selection of $106$ original panoramic images after manual screening by five examiners, thereby constructing the QuadOoS dataset. Detailed information is provided in Fig. 11(b), and see Appendix B-A for details of the types of outliers.
((a))
Frames Resolution Locations Objects Annotations
Test 106 $2048{\times}400$ 8 $\geq$18 735
((b))
((c))
Figure 12: Examples from the established QuadOoS dataset. The images come with challenges such as uneven exposure, color inconsistencies, and increased motion blur. A large number of small outliers are scattered in the background.
V Experiment Results and Analysis
V-A Datasets
@@
V-B Implementation Details
Closed-Set Training. We follow [29] and adopt the same training recipe and losses without any special design. The difference is that we adopt the ConvNeXt-Large CLIP [75] from OpenCLIP [84], pre-trained on the LAION-2B dataset [85], and freeze its weights. In reference to previous research works [28, 27], only one decoder layer with $100$ queries is used in the transformer decoder. Both dimensions $C_{m}$ and $C_{q}$ are set to $256$. The model is trained for $90k$ iterations using a batch size of $16$ on $4$ NVIDIA GeForce RTX 3090 GPUs. The training is optimized with the AdamW optimizer and weight decay $0.05$. We employ the initial learning rate $1{\times}10^{-4}$ and a polynomial schedule.
@@ V-B Implementation Details
Method DenseOoS QuadOoS Mean
PEBAL [32] 16.39 7.59 46.29 11.34 39.25 13.87 23.42
DenseHybrid [33] 24.44 7.74 45.19 19.09 33.86 21.77 20.80
EAM [26] 17.97 6.46 61.58 33.05 79.66 25.51 43.06
Mask2Anomaly [28] 36.28 8.79 48.10 25.93 92.55 31.10 50.67
RbA [27] 51.31 21.87 57.54 34.12 90.34 42.71 56.11
Ours 85.56 0.45 63.30 48.13 85.22 66.85 42.84
Method mIoU
DAFormer [34] 44.53
Trans4PASS (Tiny) [12] 44.93
Trans4PASS (Small) [12] 49.35
Trans4PASS+ (Tiny) [2] 49.29
Trans4PASS+ (Small) [2] 50.11
360SFUDA++ [48] 50.46
Ours 63.30
V-C Evaluation Metrics
@@
V-D Results of Panoramic Out-of-distribution Segmentation
Figure 13: Visual comparisons of PanOoS. We observe that when existing pinhole-OoS methods are transferred to PanOoS, current state-of-the-art methods generally struggle with misidentifying complex background as outliers. Although EAM [26] has a low false positive rate, it incorrectly identifies bicycles as outliers and nearly overlooks the actual outliers. In contrast, Mask2Anomaly [28] and RbA [27] suffer from large false positives in the background region, whereas POS delivers more precise PanOoS across the $360^{\circ}$. Zoom in for a better view.
@@ V-D Results of Panoramic Out-of-distribution Segmentation
In Fig. 13, we visualize the PanOoS results produced by POS and other mask-transformer-based methods. Notably, POS effectively suppresses false positives in background regions, especially at the boundaries separating inliers, while better preserving the smoothness of the outlier map compared to other methods, despite sharing the same mask classification training paradigm. Benefiting from the generalization ability of vision-language models, Prompt-based Restoration Attention (PRA) and Bilevel Prompt Distribution Learning (BPDL), POS performs excellently on the two benchmarks for PanOoS, perfectly adapting to the background clutter and pixel distortions caused by the broader FoV of panoramic images.
V-E Results of Pinhole Out-of-distribution Segmentation